OpenStudy (anonymous):

F=9/5C+32, solve for C

OpenStudy (dumbcow):

Subtract 32 on both sides Multiply by 5/9 on both sides

OpenStudy (anonymous):

\[C=\frac{5}{9}(F-32)\]

OpenStudy (saifoo.khan):

F - 32 = 9C/5 5(F - 32) = 9C 9C = 5(F - 32) C = 5(F - 32) -------- 9

OpenStudy (aravindg):

160C+9/5F=C